Find the side of the regular octagon inscribed in a circle of radius 50 mm. Solve in two ways. Also, find the value of each interior and exterior angles of the regular polygon.


1
Expert's answer
2021-01-06T19:55:36-0500
"\alpha=\dfrac{360\degree}{8}=45\degree"


Consider the isosceles triangle "A_1OA_2." The Law of Cosines


"a^2=R^2+R^2-2R\cdot R\cos\alpha"

"a=R\sqrt{2-\sqrt{2}}"

"a=50\sqrt{2-\sqrt{2}}\ mm"

Let "A_1(\dfrac{50\sqrt{2}}{2}, \dfrac{50\sqrt{2}}{2}), A_2(50,0)." Then


"a=A_1A_2=\sqrt{(50-\dfrac{50\sqrt{2}}{2})^2+(0-\dfrac{50\sqrt{2}}{2})^2}"

"=50\sqrt{1-\sqrt{2}+\dfrac{1}{2}+\dfrac{1}{2}}=50\sqrt{2-\sqrt{2}}\ (mm)"

"a=50\sqrt{2-\sqrt{2}}\ mm"



"\beta=\dfrac{180\degree(8-1)}{8}=157.5\degree"




"interior\ angle=157.5\degree"

The interior and exterior angle add up to 180°


"exterior\ angle=180\degree-\beta=180\degree-157.5\degree"

"exterior\ angle=22.5\degree"